OpenStudy (anonymous):

Solve in Point-Slope form: (12,4), m=0

OpenStudy (anonymous):

Alright so what's the point - slope form? :P

OpenStudy (anonymous):

y-4=0(x-12) y=4

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

\[y-4=0(x-12)\]

OpenStudy (anonymous):

\[y-4=0 => y = 4\]

OpenStudy (texaschic101):

y - 4 = 0(x - 12) y - 4 = 0 y = 4

OpenStudy (anonymous):

Thank you, i was just making sure i did the problem right!